OpenStudy (anonymous):

What is the value of z for the equation 1/2z = −3/4 + 1/4z −3 −1 1 3

OpenStudy (anonymous):

@paki

OpenStudy (perl):

is this the question: $$ \Large \rm \frac12 z = \frac{−3}{4} + \frac{1}{4}z $$

OpenStudy (anonymous):

\[-\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

with a - sign infront

OpenStudy (perl):

like this? $$ \Large \rm \frac12 z = -\frac{3}{4} + \frac{1}{4}z $$

OpenStudy (anonymous):

yea

OpenStudy (perl):

ok multiply both sides by 4

OpenStudy (anonymous):

2z=-3/4+1?

OpenStudy (perl):

almost

OpenStudy (perl):

$$ \Large \rm { \frac12 z = -\frac{3}{4} + \frac{1}{4}z \\ \color{red}4\cdot \frac12 z =\color{red}4 \cdot \left( \frac{−3}{4} + \frac{1}{4}z \right) \\ \frac42 z =\color{red}4 \cdot \frac{−3}{4} + \color{red}4 \cdot \frac{1}{4}z \\ 2 z = \cancel {\color{red}4} \cdot \frac{−3}{\cancel4} + \cancel{\color{red}4} \cdot \frac{1}{\cancel4}z } $$

OpenStudy (anonymous):

so 2z=-3+1z

OpenStudy (perl):

correct

OpenStudy (perl):

now solve for z

OpenStudy (anonymous):

x=-6?

OpenStudy (anonymous):

−3 −1 1 3 those are the choices